Two protons in an atom are 1.75 m x 10 ^ -8 m from each other. Calculate


the electric force between them.

1
Expert's answer
2022-03-09T12:19:27-0500

The Coulomb's law says

"F=k\frac{|q_1||q_2|}{r^2}"

"F=9*10^9*\frac{1.6 *10 ^ {-19}*1.6 *10 ^ {-19}}{(1.75 *10 ^ {-8})^2}=7.5*10^{-13}\:\rm N"